# Omni-Thinker: Scaling Cross-Domain Generalization in LLMs via Multi-Task RL with Hybrid Rewards

### v2

@@ A.2 Results
Figure 5: Backward transfer (BWT) matrix. Each cell shows the BWT for a target task (columns) when training a single-task GRPO model on a source task (rows). Positive BWT indicates forgetting; negative indicates positive transfer.
Table 1: Ablation study comparing curriculum Learning (Curr) to its reversed task ordering (Reverse-Curr). Ordering tasks from structured to open-ended domains consistently improves performance across all, while reversing this order leads to degraded results, especially in open-ended domains.
Task Base Model Curr Reverse-Curr
Math 55.9 59.6 58.2
General QA 41.5 52.2 22.6
Code Generation 29.6 35.4 32.7
Creative Writing 74.2 84.2 75.0
Table 2: Performance across benchmarks. MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math). Model Merge applies TIES-merging across tasks. Bolded values mark the best per row. Domains include Math (7 sets), MMLU-Pro (9 categories), Coding (2 sets), and Creative Writing (MT-Bench).
Task Base Model ST: Coding ST: Math ST: QA ST: Writing SFT Model Merging Joint MT Curriculum MT
Math
aime24 15.33 15.33 14.67 14.00 13.33 16.67 14.67 19.33 14.67
amc23 62.50 60.00 60.00 62.50 59.00 62.00 65.50 64.00 65.50
gaokao2023en 72.99 74.29 76.10 74.03 75.58 74.29 74.81 76.62 77.14
math500 78.20 78.80 80.40 75.40 79.20 76.80 79.80 77.60 81.00
minerva_math 64.34 63.97 66.54 63.24 61.76 65.07 66.18 68.38 71.69
OlympiadBench 42.07 42.96 43.70 41.33 42.96 42.96 41.78 43.56 47.41
General QA
Biology 57.60 56.76 52.30 67.36 59.00 66.25 65.55 67.22 68.76
Business 33.46 39.04 25.60 58.68 32.95 48.16 59.82 49.81 47.53
Chemistry 35.78 31.80 27.30 47.70 38.34 44.08 42.49 42.05 50.71
Computer Science 53.66 48.05 50.24 55.12 51.95 53.66 53.90 58.78 59.27
Economics 42.65 49.17 38.74 62.91 44.91 59.60 61.97 56.75 62.09
Engineering 28.28 31.27 20.43 37.46 26.63 37.77 38.08 35.81 37.05
Health 46.70 46.21 45.23 50.98 47.19 45.72 52.69 50.73 57.09
History 37.27 33.33 34.65 47.24 38.58 33.86 47.24 43.31 45.67
Law 23.16 23.98 20.62 27.88 23.25 26.79 26.61 27.52 29.70
Math 55.37 52.63 50.41 59.29 56.25 57.36 58.25 59.22 61.21
Other 44.26 40.04 39.72 50.97 43.94 46.43 51.84 49.89 53.25
Philosophy 36.87 34.27 33.27 43.89 35.47 38.20 41.48 42.08 42.89
Physics 41.11 37.41 30.79 53.66 41.57 49.81 46.73 48.04 55.58
Psychology 50.88 51.50 45.36 60.15 51.75 59.02 59.40 59.27 61.78
Code Generation
BigCodeBench 46.49 50.35 46.66 47.10 46.84 44.47 48.07 47.19 49.47
LiveCodeBench 12.71 21.80 13.07 13.79 13.31 14.21 20.78 17.04 21.30
Creative Writing
MT-Bench (Writing) 74.16 71.60 74.16 63.00 78.33 44.17 67.50 83.00 84.17
@@ A.3 Detailed Hyper-Parameters
We show the hyperparameters used in our training in Table3
Table 3: Training Hyperparameters for All Training Settings. MT = Multi-Task RL, ST = Single-Task RL (e.g., ST: Math = RL trained only on math).
Hyperparameter Curr. MT Joint MT ST:Coding ST:Math ST:QA ST:Writing SFT
@@
   Train Batch Size 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 256$\times$6 128

### v3

@@ A.2 Results
A.2.1 Output Format Matters: Full-Text Answers Enhance Generalization
We examine how output format impacts generalization by comparing models trained to generate full-text answers versus selecting letter choices in multiple-choice QA (MCQ). Using GRPO, we train two single-task policies on the training set, one prompted to produce full-text final answers at the end of its chain-of-thought completions, and the other to output only letter choices (e.g., “A”, “B”, “C”).
As shown in Figure 3, the model trained to output full-text answers achieves significantly better generalization when evaluated with free-form QA prompts on MMLU Pro (51% vs. 41%). While the letter-choice model slightly outperforms when evaluated strictly on MCQ prompts, the full-text model remains competitive across both prompt formats.
A.2.2 Curriculum Accuracies Predictions
On Figure 4 are provided BWT matrices for validation sets and test sets. We compute the ranking of curricula using both, see table 4. We observe that the Forgettability heuristic provides the second best prediction on test BWT, but exact LOM yields a potentially better candidate. Using validation BWT, the curriculum Code-Math-QA-Writing comes second,, but the proposed exact LOM should be rejected based on our entropy discussion.
Figure 4: Comparison of backward transfer on validation set (left) and test set (right).
Table 4: Curricula ranked by final geometrical average relative improvement $\overline{\Delta_{\mathrm{rel}}}:=\frac{1}{K}\log{\prod_{k=1}^{K}\frac{\mathrm{Acc}(\theta_{s_{\mathrm{max}}},T_{k})}{\mathrm{Acc}(\theta_{0},T_{k})}}$ on test set (left block) and eval set (right block), with predicted accuracies (%) for each task at each order, and their ranking scores.
Test set order Validation set order
Order M C Q W $\overline{\Delta_{\mathrm{rel}}}$ Order M C Q W $\overline{\Delta_{\mathrm{rel}}}$
$\mathcal{M}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{W}$ 55.2 37.7 51.9 78.4 12.85 $\mathcal{C}\,\mathcal{W}\,\mathcal{M}\,\mathcal{Q}$ 34.9 42.5 26.8 35.4 26.00
$\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ 57.3 35.4 51.9 78.4 12.22 $\mathcal{C}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{W}$ 36.7 42.5 27.5 32.6 25.85
$\mathcal{M}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{W}$ 55.2 36.7 51.0 78.4 11.70 $\mathcal{C}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{M}$ 33.1 42.5 28.1 35.3 25.81
$\mathcal{M}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{C}$ 55.2 36.1 51.0 75.7 10.43 $\mathcal{W}\,\mathcal{C}\,\mathcal{M}\,\mathcal{Q}$ 34.9 44.7 26.8 33.4 25.77
$\mathcal{M}\,\mathcal{C}\,\mathcal{W}\,\mathcal{Q}$ 55.2 37.7 51.3 66.6 8.49 $\mathcal{C}\,\mathcal{M}\,\mathcal{W}\,\mathcal{Q}$ 36.7 42.5 26.8 32.8 25.29
$\mathcal{C}\,\mathcal{M}\,\mathcal{W}\,\mathcal{Q}$ 57.3 35.4 51.3 66.6 7.86 $\mathcal{C}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{M}$ 33.1 42.5 27.4 35.4 25.25
$\mathcal{M}\,\mathcal{W}\,\mathcal{C}\,\mathcal{Q}$ 55.2 37.1 51.3 64.3 7.22 $\mathcal{C}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{W}$ 34.8 42.5 28.1 32.6 25.09
$\mathcal{C}\,\mathcal{W}\,\mathcal{M}\,\mathcal{Q}$ 57.2 35.4 51.3 64.3 6.96 $\mathcal{W}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{M}$ 33.1 44.7 27.4 33.4 25.02
$\mathcal{W}\,\mathcal{M}\,\mathcal{C}\,\mathcal{Q}$ 55.1 37.1 51.3 62.1 6.32 $\mathcal{M}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{W}$ 37.0 40.2 27.5 32.6 24.71
$\mathcal{M}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{C}$ 55.2 36.1 50.4 64.3 6.07 $\mathcal{W}\,\mathcal{M}\,\mathcal{C}\,\mathcal{Q}$ 35.3 42.3 26.8 33.4 24.64
$\mathcal{W}\,\mathcal{C}\,\mathcal{M}\,\mathcal{Q}$ 57.2 34.8 51.3 62.1 5.69 $\mathcal{M}\,\mathcal{C}\,\mathcal{W}\,\mathcal{Q}$ 37.0 40.2 26.8 32.8 24.16
$\mathcal{C}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{W}$ 57.7 35.4 39.5 78.4 5.56 $\mathcal{M}\,\mathcal{W}\,\mathcal{C}\,\mathcal{Q}$ 37.0 42.3 26.8 30.9 23.93
$\mathcal{W}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{C}$ 55.1 36.1 50.4 62.1 5.18 $\mathcal{Q}\,\mathcal{C}\,\mathcal{W}\,\mathcal{M}$ 33.1 41.3 26.1 35.3 23.23
$\mathcal{Q}\,\mathcal{M}\,\mathcal{C}\,\mathcal{W}$ 55.6 36.7 38.8 78.4 5.05 $\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ 33.1 43.4 26.1 33.3 23.00
$\mathcal{C}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{M}$ 57.6 35.4 39.5 75.7 4.67 $\mathcal{Q}\,\mathcal{C}\,\mathcal{M}\,\mathcal{W}$ 34.8 41.3 26.1 32.6 22.52
$\mathcal{Q}\,\mathcal{C}\,\mathcal{M}\,\mathcal{W}$ 57.7 34.4 38.8 78.4 4.41 $\mathcal{W}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{M}$ 33.1 43.4 25.4 33.4 22.45
$\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ 55.6 36.1 38.8 75.7 3.77 $\mathcal{M}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{W}$ 37.0 39.1 25.5 32.6 22.14
$\mathcal{Q}\,\mathcal{C}\,\mathcal{W}\,\mathcal{M}$ 57.6 34.4 38.8 75.7 3.52 $\mathcal{W}\,\mathcal{M}\,\mathcal{Q}\,\mathcal{C}$ 35.3 41.1 24.9 33.4 22.07
$\mathcal{Q}\,\mathcal{W}\,\mathcal{M}\,\mathcal{C}$ 55.5 36.1 38.8 73.1 2.88 $\mathcal{M}\,\mathcal{Q}\,\mathcal{W}\,\mathcal{C}$ 37.0 41.1 25.5 30.8 21.91
$\mathcal{Q}\,\mathcal{W}\,\mathcal{C}\,\mathcal{M}$ 57.6 33.9 38.8 73.1 2.25 $\mathcal{Q}\,\mathcal{W}\,\mathcal{M}\,\mathcal{C}$ 33.5 41.1 26.1 33.3 21.87
$\mathcal{C}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{M}$ 57.6 35.4 39.0 64.3 0.31 $\mathcal{Q}\,\mathcal{M}\,\mathcal{C}\,\mathcal{W}$ 35.2 39.1 26.1 32.6 21.39
$\mathcal{W}\,\mathcal{C}\,\mathcal{Q}\,\mathcal{M}$ 57.6 34.8 39.0 62.1 -0.97 $\mathcal{M}\,\mathcal{W}\,\mathcal{Q}\,\mathcal{C}$ 37.0 41.1 24.9 30.9 21.35
$\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ 55.5 36.1 38.4 62.1 -1.48 $\mathcal{W}\,\mathcal{Q}\,\mathcal{M}\,\mathcal{C}$ 33.5 41.1 25.4 33.4 21.31
$\mathcal{W}\,\mathcal{Q}\,\mathcal{C}\,\mathcal{M}$ 57.6 33.9 38.4 62.1 -2.11 $\mathcal{Q}\,\mathcal{M}\,\mathcal{W}\,\mathcal{C}$ 35.2 41.1 26.1 30.8 21.16
@@ A.3 Detailed Hyper-Parameters
We summarize the hyperparameters used in our experiments in Table 5. These values were chosen through a combination of prior work, small-scale ablations, and practical compute considerations.
Table 5: Training Hyperparameters for All Training Settings. ST = Single-Task RL (e.g., ST Math = RL trained only on math).
Hyperparameter Curr. Learning Joint Training ST Coding ST Math ST QA ST Writing SFT
@@
   Train Batch Size $\times6 25$$\times6$ $\times6 25$$\times6$ $\times6 25$$\times6$ 128